What is the probability of cat weight between 5.25 kg and 6.81 kg


1
Expert's answer
2021-12-15T08:38:00-0500

Let "X=" the cat weight: "X\sim N(\mu, \sigma^2)."

Given "\mu=4.76kg, \sigma=1.23kg."


"P(5.25<X<6.81)=P(X<6.81)-P(X\leq 5.25)"

"=P(Z<\dfrac{6.81-4.76}{1.23})-P(Z\leq \dfrac{5.25-4.76}{1.23})"

"\approx P(Z<1.666667)-P(Z\leq0.398374 )"

"\approx 0.95221-0.65482\approx0.2974"

The probability of cat weight between 5.25 kg and 6.81 kg is 0.2974.